# Effective Interplay between Sparsity and Quantization: From Theory to Practice

### v1

@@ 6 Conclusion
We provide a comprehensive analysis of the interplay between quantization and sparsity in DNNs, showing that applying sparsity before quantization (SQ) minimizes additional errors and yields better model accuracy. Moreover, our mathematical analysis and extensive empirical study with large language models (OPT, LLaMA) and vision transformers (ViT) demonstrate that quantization and sparsity are not orthogonal and their combined use can adversely affect model accuracy. Our findings provide valuable insights for optimizing the compression of large models while preserving accuracy.
Acknowledgments and Disclosure of Funding
@@ Appendix A Additional related work
Quantization. Element-wise scaling formats, such as FP32, BFLOAT32 [60], and FP16 [38], consist of sign, mantissa, and exponent components, differing in the bit allocation for each component. Conversely, block-wise scaling formats assign scaling factors to blocks of elements, with block sizes varying by format. For instance, INT8 employs per-tensor scaling, where a single scaling factor is shared by around 1K elements.
Recent research highlights the effectiveness of fine-grained block-wise scaling formats with block sizes smaller than 100 elements, especially in the sub-8-bit regime for both training and inference [11, 49, 50, 70, 6]. These formats are further categorized into single-level and two-level scaling. Single-level block-wise formats, such as HBFP [11, 6] and MXINT [50], enable fixed-point arithmetic by sharing a single exponent within a block of mantissa or integers. Two-level formats, like MXFP [49, 50, 40] and FP8, use more granular scaling factors at the second level, offering greater robustness across diverse range of models.
Sparsity. Unstructured sparsity [18, 17, 13, 12] involves removing individual tensor elements without any specific pattern. Structured sparsity [62], on the other hand, employs specific patterns when pruning tensor elements. Recent work [65, 26] has highlighted the effectiveness of fine-grained N:M structured sparsity in mitigating model accuracy loss. The introduction of the 2:4 structured-sparse Tensor Core in the Nvidia Ampere architecture [44] has further driven research in developing N:M sparsity training recipes [42, 48, 71, 54, 24, 34]. The fundamental operation in any sparsification scheme is selecting candidate elements for pruning among which magnitude-based sparsity [19] is one of the most widely used methods [34, 9, 2]. In addition, recent work has introduced one-shot pruning methods, such as SparseGPT [15] and Wanda [53], aiming to eliminate the need for an additional fine-tuning phase. While these methods achieve state-of-the-art performance for one-shot pruning, evidence suggests that incorporating a fine-tuning phase can lead to better model quality [53, 55]. Although these methods are proposed to eliminate fine-tuning and present state-of-the-art accuracies for one-shot pruning, it has been shown that fine-tuning still improves accuracies significantly [53, 55].
@@ Appendix D Fine-tuning strategies
Appendix E Post-training one-shot sparsity methods
@@
50% | Magnitude | S Q | 29.94 | 30.22 | 31.13 | 31.20 | 30.46 | 32.51 | 6.31 | 6.94 | 6.40 | 6.38 | 6.32 | 6.51 |
Q S | - | 34.71 | 36.39 | 35.60 | 37.48 | 40.86 | - | 8.13 | 8.47 | 9.32 | 9.86 | 10.20 | ||
Wanda | S Q | 38.97 | 39.29 | 39.72 | 40.02 | 39.21 | 42.33 | 6.46 | 6.47 | 6.53 | 6.53 | 6.48 | 6.73 | |
Q S | - | 40.01 | 40.58 | 40.43 | 40.52 | 42.57 | - | 6.46 | 6.55 | 6.52 | 6.48 | 6.79 | ||
SparseGPT | S Q | 33.24 | 33.22 | 35.27 | 34.22 | 33.41 | 35.86 | 6.51 | 6.51 | 6.58 | 6.58 | 6.52 | 6.77 | |
Q S | - | 33.54 | 35.32 | 34.29 | 33.64 | 36.80 | - | 6.53 | 6.60 | 6.58 | 6.55 | 6.93 | ||
2:4 | Magnitude | S Q | 31.89 | 32.76 | 33.99 | 33.41 | 32.25 | 34.58 | 9.30 | 9.37 | 9.35 | 9.32 | 9.39 | 10.68 |
Q S | - | 45.06 | 44.16 | 42.25 | 46.57 | 55.64 | - | 14.65 | 14.35 | 14.50 | 14.98 | 18.64 | ||
Wanda | S Q | 79.91 | 79.81 | 85.25 | 84.10 | 80.62 | 90.66 | 11.36 | 11.37 | 11.15 | 11.35 | 11.45 | 12.74 | |
Q S | - | 80.28 | 86.69 | 84.38 | 80.69 | 91.04 | - | 11.28 | 11.24 | 11.46 | 11.36 | 13.61 | ||
SparseGPT | S Q | 45.14 | 45.34 | 48.44 | 46.49 | 45.52 | 50.74 | 10.22 | 10.21 | 10.15 | 10.26 | 10.26 | 10.86 | |
Q S | - | 44.96 | 48.67 | 46.50 | 45.82 | 57.39 | - | 10.21 | 10.24 | 10.26 | 10.21 | 11.16 |
@@
We further explore the effectiveness of post-training one-shot sparsity methods, specifically SparseGPT and Wanda, which utilize a selection criterion based on the product of the magnitudes of weights and activations. We report our results in Table 6 We observe that magnitude-based sparsity continues to achieve better perplexities due to fine-tuning. However, because of their selection criterion, SparseGPT and Wanda are not affected by the order of the operations. Even when quantization alters the relative magnitudes within a weight tensor, the corresponding activations can compensate by preserving the original ranking of importance when multiplied together. Consequently, the difference in perplexities between SQ and QS for these methods is minimal, and in few instances, QS yields better perplexities.
Appendix F Orthogonality bound for ViT
@@
Table 7 shows the results of cross-entropy loss across various combinations of sparsity and quantization schemes for ViT-B/16 performing the image classification task on ImageNet-1k. First, we note that the calculated orthogonality bound serves as a correct lower bound for most configurations, supporting our theoretical analysis. Second, ViT-B/16 is significantly more robust to the combination of quantization and sparsity schemes compared to the other LLMs studied in this paper. When used with moderate sparsity levels (50% and 2:4) and 8-bit/6-bit number formats, the actual cross-entropy loss is close to the calculated orthogonality bound, showing the robustness of ViT-B/16. Only at higher compression rates achieved by using 75% or 1:4 sparsity and 4-bit number formats such as HBFP4, do we see the impact of the quantization and sparsity errors affecting the final cross-entropy loss, making it significantly higher than the calculated orthogonality bound.
@@
Metric | Orthogonality Bound | ||||||
@@
Appendix G Layer-wise error analysis
Our mathematical study focuses on the per-layer error caused by the combination of quantization and sparsity. To demonstrate that this error is additive and increases with the layer index, we inspect the outputs of intermediate layers. In our experiment, we apply quantization and magnitude-based sparsity to all linear layers of the pre-trained OPT-125M model in a zero-shot manner. We feed a sample from a test subset to the compressed model and the corresponding full precision dense model and measure the difference between Feed-Forward outputs for each Transformer block. We repeat this experiment for SQ and QS orders of transformations and show comparison results in Figure 1.
We observe that the per-layer error increases monotonically with layer index, regardless of the order of transformations, and reaches its maximum value after the final layer. This error directly affects the output of the model and results in the metric degradation, which we have detected in the previous experiments 3.1, 3.2.
Layer-wise analysis also supports our theoretical findings that the optimal order of transformations is sparsity followed by quantization. The SQ order of transformations consistently yields lower errors in the intermediate layer outputs compared to the reverse order. We observe an exponential increase in error with the QS order of transformations for the final layers, resulting in an error almost four times larger than with the optimal order.
@@ Appendix H Robustness analysis
Appendix I Proofs for the mathematical analysis
@@
There exists a permutation of the vector such that maps the element in to the element in . This helps us to get an optimal upper bound for the additional error:
For the case of sparsity then number of nullified elements within the block equals . Therefore, with respect to norm, the last error term can be evaluated as follows:
@@
Appendix J Definitions of for numerical formats
Appendix K Analysis of the upper bound of the dot product error
K.1 Upper bound is reachable
@@
K.2 Contribution of additional error terms
@@
Figures 4 and 5 depict the results of the experiment. If we consider the order in Figure 4, we can see that the term advocates for almost half of the error of the composition. However, in the order the term has a much lower impact, which proves our first hypothesis.

### v2

@@ 6 Conclusion
We provide a comprehensive analysis of the interplay between sparsity and quantization in DNNs, showing that applying sparsity before quantization (SQ) minimizes additional errors and yields better model accuracy. Moreover, our mathematical analysis and extensive empirical study with large language models (OPT, LLaMA), vision transformers (ViT), and convolutional neural networks (ResNet) demonstrate that sparsity and quantization are non-orthogonal and their combined use can adversely affect model accuracy. Our findings provide valuable insights for optimizing the compression of large models while preserving accuracy.
Acknowledgments
@@ Appendix A Additional related work
Quantization. Element-wise scaling formats, such as FP32, BFLOAT32 (Wang & Kanwar, 2019), and FP16 (Micikevicius et al., 2018), consist of sign, mantissa, and exponent components, differing in the bit allocation for each component. Conversely, block-wise scaling formats assign scaling factors to blocks of elements, with block sizes varying by format. For instance, INT8 employs per-tensor scaling, where a single scaling factor is shared by around 1K elements.
Recent research highlights the effectiveness of fine-grained block-wise scaling formats with block sizes smaller than 100 elements, especially in the sub-8-bit regime for both training and inference (Drumond et al., 2018a; Rouhani et al., 2023a; b; Zhang et al., 2022a; Darvish Rouhani et al., 2020a). These formats are further categorized into single-level and two-level scaling. Single-level block-wise formats, such as HBFP (Drumond et al., 2018a; Darvish Rouhani et al., 2020a) and MXINT (Rouhani et al., 2023b), enable fixed-point arithmetic by sharing a single exponent within a block of mantissa or integers. Two-level formats, like MXFP (Rouhani et al., 2023a; b; Micikevicius et al., 2022) and FP8, use more granular scaling factors at the second level, offering greater robustness across diverse range of models.
Sparsity. Unstructured sparsity (Han et al., 2015a; Guo et al., 2016; Frankle & Carbin, 2019; Evci et al., 2020a) involves removing individual tensor elements without any specific pattern. Structured sparsity (Wen et al., 2016), on the other hand, employs specific patterns when pruning tensor elements. Recent work (Yao et al., 2019; Kang, 2020) has highlighted the effectiveness of fine-grained N:M structured sparsity in mitigating model accuracy loss. The introduction of the 2:4 structured-sparse Tensor Core in the Nvidia Ampere architecture (Nvidia, 2021) has further driven research in developing N:M sparsity training recipes (Mishra et al., 2021b; Pool & Yu, 2021; Zhou et al., 2021; Sun et al., 2021; Hubara et al., 2021; Lu et al., 2023).
The fundamental operation in any sparsification scheme is selecting candidate elements for pruning among which magnitude-based sparsity (Han et al., 2015b) is one of the most widely used methods (Lu et al., 2023; Ding et al., 2023; Bambhaniya et al., 2024). In addition, recent work has introduced one-shot pruning methods, such as SparseGPT (Frantar & Alistarh, 2023) and Wanda (Sun et al., 2023), aiming to eliminate the need for an additional fine-tuning phase. While these methods achieve state-of-the-art performance for one-shot pruning, evidence suggests that incorporating a fine-tuning phase can lead to better model quality (Sun et al., 2023; Syed et al., 2023). Although these methods are proposed to eliminate fine-tuning and present state-of-the-art accuracies for one-shot pruning, it has been shown that fine-tuning still improves accuracies significantly (Sun et al., 2023; Syed et al., 2023).
@@ Appendix D Fine-tuning strategies
Appendix E Experimental Setup Summary
Table 6 presents a summary of our experimental setup.
Compressed operations by model type | |||||
LLMs | ViTs | CNNs | |||
Linear Layers | Yes | Yes | Yes | ||
Convolutional Layers | – | – | Yes | ||
Embedding Layer | No | Yes | – | ||
Attention MatMuls | Yes | Yes | – | ||
Tensor-level granularity for compressed operations | |||||
Weights | Activations | Gradients | |||
Forward Pass (for inference and finetuning) | Sparsified + quantized | Quantized | – | ||
Backward Pass (for finetuning only) | Quantized | Quantized | Quantized | ||
(1) Note: Only matrix multiplications are compressed; other operations (e.g., optimizer updates) are in FP32 | |||||
(2) Note: Master weights are stored in FP32, since our primary aim is to compress inference. | |||||
Workflow for the two different compression orders | |||||
S Q |
Q S |
Appendix F Post-training one-shot sparsity methods
@@
50% | Magnitude | S Q | 29.94(+2.29) | 30.22(+2.16) | 31.13(+2.68) | 31.20(+3.19) | 30.46(+2.65) | 32.51(+2.60) | 6.31(+1.19) | 6.94(+1.79) | 6.40(+1.23) | 6.38(+1.22) | 6.32(+1.2) | 6.51(+1.27) |
Q S | - | 34.71(+6.65) | 36.39(+7.94) | 35.60(+7.59) | 37.48(+9.67) | 40.86(+10.95) | - | 8.13(+2.98) | 8.47(+3.30) | 9.32(+4.16) | 9.86(+4.74) | 10.20(+4.96) | ||
Wanda | S Q | 38.97(+11.32) | 39.29(+11.23) | 39.72(+11.27) | 40.02(+12.01) | 39.21(+11.40) | 42.33(+12.42) | 6.46(+1.34) | 6.47(+1.32) | 6.53(+1.36) | 6.53(+1.37) | 6.48(+1.36) | 6.73(+1.49) | |
Q S | - | 40.01(+11.95) | 40.58(+12.13) | 40.43(+12.42) | 40.52(+12.71) | 42.57(+12.66) | - | 6.46(+1.31) | 6.55(+1.38) | 6.52(+1.36) | 6.48(+1.36) | 6.79(+1.55) | ||
SparseGPT | S Q | 33.24(+5.59) | 33.22(+5.16) | 35.27(+6.82) | 34.22(+6.21) | 33.41(+5.60) | 35.86(+5.95) | 6.51(+1.39) | 6.51(+1.36) | 6.58(+1.41) | 6.58(+1.42) | 6.52(+1.40) | 6.77(+1.53) | |
Q S | - | 33.54(+5.48) | 35.32(+6.87) | 34.29(+6.28) | 33.64(+5.83) | 36.80(+6.89) | - | 6.53(+1.38) | 6.60(+1.43) | 6.58(+1.42) | 6.55(+1.43) | 6.93(+1.69) | ||
2:4 | Magnitude | S Q | 31.89(+4.24) | 32.76(+4.7) | 33.99(+5.54) | 33.41(+5.40) | 32.25(+4.44) | 34.58(+4.67) | 9.30(+4.18) | 9.37(+4.22) | 9.35(+4.18) | 9.32(+4.16) | 9.39(+4.27) | 10.68(+5.44) |
Q S | - | 45.06(+17.00) | 44.16(+15.71) | 42.25(+14.24) | 46.57(+18.76) | 55.64(+25.73) | - | 14.65(+9.50) | 14.35(+9.18) | 14.50(+9.34) | 14.98(+9.86) | 18.64(+13.40) | ||
Wanda | S Q | 79.91(+52.26) | 79.81(+51.75) | 85.25(+56.80) | 84.10(+56.09) | 80.62(+52.81) | 90.66(+60.75) | 11.36(+6.24) | 11.37(+6.22) | 11.15(+5.98) | 11.35(+6.19) | 11.45 (+6.33) | 12.74 (+7.50) | |
Q S | - | 80.28(+52.22) | 86.69(+58.24) | 84.38(+56.37) | 80.69(+52.88) | 91.04(+61.13) | - | 11.28(+6.13) | 11.24(+6.07) | 11.46(+6.30) | 11.36(+6.24) | 13.61(+8.37) | ||
SparseGPT | S Q | 45.14(+17.49) | 45.34(+17.28) | 48.44(+19.99) | 46.49(+18.48) | 45.52(+17.71) | 50.74(+20.83) | 10.22(+5.10) | 10.21(+5.06) | 10.15(+4.98) | 10.26(+5.10) | 10.26(+5.14) | 10.86(+5.62) | |
Q S | - | 44.96(+16.9) | 48.67(+20.22) | 46.50(+18.49) | 45.82(+18.01) | 57.39(+27.48) | - | 10.21(+5.06) | 10.24(+5.07) | 10.26(+5.10) | 10.21(+5.09) | 11.16(+5.92) |
@@
We further explore the effectiveness of post-training one-shot sparsity methods, specifically SparseGPT and Wanda, which utilize a selection criterion based on the product of the magnitudes of weights and activations. We report our results in Table 7. We observe that magnitude-based sparsity continues to achieve better perplexities due to fine-tuning. However, because of their selection criterion, SparseGPT and Wanda are not affected by the order of the operations. Even when quantization alters the relative magnitudes within a weight tensor, the corresponding activations can compensate by preserving the original ranking of importance when multiplied together. Consequently, the difference in perplexities between SQ and QS for these methods is minimal, and in few instances, QS yields better perplexities.
Appendix G Orthogonality threshold for ViT
@@
Table 8 shows the results of cross-entropy loss across various combinations of sparsity and quantization schemes for ViT-B/16 performing the image classification task on ImageNet-1k. First, we note that the calculated orthogonality threshold serves as a correct lower bound for most configurations, supporting our mathematical analysis. Second, ViT-B/16 is significantly more robust to the combination of sparsity and quantization schemes compared to the other LLMs studied in this paper. When used with moderate sparsity levels (50% and 2:4) and 8-bit/6-bit number formats, the actual cross-entropy loss is close to the calculated orthogonality threshold, showing the robustness of ViT-B/16. Only at higher compression rates achieved by using 75% or 1:4 sparsity and 4-bit number formats such as HBFP4, do we see the impact of the sparsity and quantization errors affecting the final cross-entropy loss, making it significantly higher than the calculated orthogonality threshold.
@@
Metric | Orthogonality Threshold | ||||||
@@ Appendix H Robustness analysis
Appendix I Layer-wise Error Propagation
We also examine the error propagation to ensure that errors introduced in earlier layers do not vanish in subsequent layers. Figure 2 illustrates how compression errors persist and propagate throughout the pre-trained OPT-2.7b model. In this analysis, we quantize each layer in isolation and measure the corresponding propagated error to the other layers, which remain in full precision. We observe that relatively large errors introduced in earlier layers, despite small fluctuations, stay on the same level as they go through the network. As a result, even a single-layer error can significantly degrade model performance, highlighting the potential threats of errors due to the non-orthogonality of compression techniques or applying them in the suboptimal order.
Appendix J Proofs for the mathematical analysis
@@ Appendix I Layer-wise Error Propagation
Let represent the number of elements pruned from the block by the sparsity transformation. Without loss of generality, we assume that the last elements in the block are pruned, as permuting the elements does not affect the block’s norm. As the sparsity transformation does not prune the largest element in the block, the parameter of quantization remains unchanged. Consequently, the quantization error for the non-zero components before and after sparsity remains the same:
For norms, where , the upper bound is attainable under one of two conditions: either the pruned elements in are originally zero, or the quantization error for all elements of is zero.
In the first case, the first inequality becomes an equality because . In the second case, the first inequality also becomes an equality because .
Similarly, the second inequality becomes an equality as quantization maps zero to zero. Thus, the quantization error for elements in is zero either because the quantization error for all elements of is zero or because the elements were originally zero: .
For norm there exist a non-trivial case. We consider the block of floating-point numbers , INT4 quantization and sparsity.
The L1-norms of the transformation errors are the following:
Therefore, is attainable. ∎
Proof of Theorem 3.6.
We consider norms, where . In these norms, . Therefore:
Thus, for this particular input , the inequality holds true. ∎
Theorem J.1 (Upper-bound of the error for suboptimal order, general case).
Let be the max-scaled block-wise quantization and be the magnitude-based N:M sparsity transformation. Let be the least upper bound for the magnitude of the quantization error for one element: . Let be a vector with ones and zeros in any order. Then the error of the composition with respect to norm has the following upper bound:
Proof of Theorem J.1.
@@
There exists a permutation of the vector such that maps the element in to the element in . Therefore, an upper bound for is:
For the case of sparsity the number of nullified elements within the block equals . Therefore:
As a result, the upper bound for the error of the composition is the following:
@@
As a corollary of Theorem J.1, with respect to norm, the last error term can be evaluated as follows:
Proof of Theorem 3.10.
Error of the composition can be written as the following:
After adding the norms, we obtain the following:
where and are the additional error terms.
@@
Theorem J.2.
Let be the max-scaled block-wise quantization and be the magnitude-based N:M sparsity transformation. Then:
Proof.
If sparsity is applied first, then from the proof of Theorem 3.5
If quantization is applied first, there are two cases.
Case 1: there are no new duplicates. Quantization cannot reorder elements, it can only make them duplicates. Therefore, if there are no new duplicates, the same elements will be sparsified after quantization as the order of the elements did not change, and the error vector will be the same:
Case 2: there are new duplicates. Let and be such elements that and , and the -th element gets sparsified instead of the -th. In this case:
Therefore,
If we consider the case and , then
Otherwise, we assume that , and have the same sign.
Here we have three subcases:
. Then
(52) -
. Then
(53) -
. Then
Therefore,
As a result,
Appendix K Definitions of for numerical formats
Appendix L Analysis of the upper bound of the dot product error
L.1 Upper bound is reachable
@@
L.2 Contribution of additional error terms
@@
Figures 5 and 6 depict the results of the experiment. If we consider the order in Figure 5, we can see that the term advocates for almost half of the error of the composition. However, in the order the term has a much lower impact, which proves our first hypothesis.